OpenStudy (anonymous):

Write the explicit formula for the geometric sequence. a1 = -5 a2 = 20 a3 = -80

OpenStudy (anonymous):

A. an = -5 • (-4)n B. an = -5(-4)n-1 C. an = -4(-5)n-1 D. an = -5 • (4)n

OpenStudy (tkhunny):

What say you? Try them out by producing those three values.

OpenStudy (anonymous):

I don't really know how to do these. I didn't get much time to study... Could you explain?

OpenStudy (tkhunny):

Use the value n = 1 in each description. See which produce -5 A. an = -5 • (-4)n a1 = -5*(-4)^1 = -5*-4 = 20 -- That's not -5 B. an = -5(-4)n-1 a1 = -5(-4)^0 = -5*1 = -4 -- Okay, this one is in the running C. an = -4(-5)n-1 a1 - -4(-5)^0 = -4*1 = -4 -- That's not -5 D. an = -5 • (4)n a1 = -5(4)^1 = -6*4 = -20 -- That's not -5 Looks like you need to try "B". What is a2? What is a3? Do we have a complete match?

OpenStudy (anonymous):

I'm still not understand.